OpenStudy (anonymous):

PLEASE HELP solve x^2-5x=3 with the quadratic formula!!

OpenStudy (zzr0ck3r):

rewrite it in the form \(ax^2+bx+c=0\) then use the quadratic formula \(x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

I know how... its just I need it and howe to solve it right away!

OpenStudy (zzr0ck3r):

look up \(\huge \uparrow\)

OpenStudy (zzr0ck3r):

there is that method and one other method, but if you don't like this one, I very much doubt you will like the other.

OpenStudy (zzr0ck3r):

"I know how.." I am confused.

OpenStudy (anonymous):

no wait let me explain. I know how to do it. Its that Im in a hurry and was wondering if you could solve it because I have 5 others to solve!

OpenStudy (zzr0ck3r):

you could have been done by now.

OpenStudy (zzr0ck3r):

I have done it twice

OpenStudy (anonymous):

alright...

OpenStudy (zzr0ck3r):

We are not here to take test or do your homework for you, we are here to guide you to an answer.

OpenStudy (anonymous):

its not homework! or a test

OpenStudy (zzr0ck3r):

Well then why do you need it done? If its not homework or a test, then there is no point in me doing something I know how to do for practice.

OpenStudy (anonymous):

Im doing it for my mom

OpenStudy (zzr0ck3r):

Well we are not here to do her homework or take her test either:)

OpenStudy (anonymous):

shhhh.. thats not what i meant

OpenStudy (zzr0ck3r):

hmm I am still confused.

OpenStudy (anonymous):

fine.. then could you help me figure it out?

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

subtract 3 from both sides and you have \(x^2-5x-3=0\) do you see that?

OpenStudy (zzr0ck3r):

\(ax^2+bx+c=0\) \(x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (zzr0ck3r):

can you tell me what a=?

OpenStudy (anonymous):

where did you get the 3 from?

OpenStudy (zzr0ck3r):

\( x^2-5x=3\)

OpenStudy (zzr0ck3r):

subtract 3 from both sides \( x^2-5x-3=3-3\) \(x^2-5x-3=0\)

OpenStudy (anonymous):

oh wait!!

OpenStudy (zzr0ck3r):

the reason we do this is to get it in the form \(ax^2+bx+c=0\) then we can use \(x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (anonymous):

\[x^{2}-5x=3\] There are 2 real solutions The Work \[x=\frac{ -b \pm \sqrt{b ^{2}- 4(a)(c)} }{ 2(a) }\] \[x=\frac{ -5 \pm \sqrt{5 ^{2}- 4(1)(3)} }{ 2(1) }\] \[\frac{ -5\pm \sqrt{13} }{ 2 }\] Solutions: x=−0.6972243622680054 x=−4.302775637731995

OpenStudy (zzr0ck3r):

why would you do that @TwinYang ? when you see someone trying to do it them selves

OpenStudy (zzr0ck3r):

If I was helping someone in a tutor lab would you just walk up and hand them the answer?

OpenStudy (zzr0ck3r):

lame

OpenStudy (anonymous):

not this problem...its x^2+4x-6=0

OpenStudy (zzr0ck3r):

see now hes tryong to cheat you out of another one.

OpenStudy (anonymous):

i already knew the answer

OpenStudy (anonymous):

IM ASKING YOU ZZR0CK3R

OpenStudy (anonymous):

fine